Find the Tangent and Normal line to the given curve. y=โˆš(16+x^2) at the origin


1
Expert's answer
2021-01-07T17:59:16-0500

Solution


  • At x=0 y=4: means that the curve does not pass through the origin.
  • The curve is a parabola having a minimum at (0,4) towards infinite y.
  • Therefore, a tangent & a normal could be drawn at (0,4) not at the origin (0,0).
  • Considering the derivative of the curve,

"\qquad\qquad\,\,\,\,\,\,\frac{dy}{dx}=\frac{1}{2\sqrt{16+x^2}}(2x)=\frac{x}{\sqrt{16+x^2}}\\
\qquad\qquad\frac{dy}{dx}_{(0,4)}=0"

  • Therefore, tangent is a line having zero gradient "\to" parallel to the x axis"\to" y= 4
  • Then the normal line is parallel to the y axis"\to" x=0"\to" the y axis.


  • Tangent: y=4
  • Normal: x=0